OpenStudy (anonymous):

Which of the following is equal to 4.5 mm^3? A) 4.5 x 10^-6 m^3 B) 4.5 x 10^-9 m^3 C) 4.5 x 10^-6 cm^3 D) 4.5 x 10^3 cm^3 Please explain how you found the answer. :)

OpenStudy (rulnick):

B

OpenStudy (rulnick):

... mm = 10^-3 m, so mm^3 = 10^-9 m^3.

OpenStudy (anonymous):

as @rulnick said \[\LARGE 4.5mm^3=4.5(m\cdot 10^{-3})^3=4.5\cdot 10^{-9} m^3\]

OpenStudy (anonymous):

Thank you :)